Your job is to lift 25.0 kg packaged baxes a vertical distance of 1.0 m from the ground onto the bed of a delivery truck. How many packaged boxes would you have to load onto the truck in one minute for your average output power that goes into lifting the boxes to be 0.11 hp?

1
Expert's answer
2022-04-19T10:00:33-0400
"P=\frac{W}{t}=\frac{Nmgh}{t}""N=\frac{Pt}{mgh}\\
=\rm\frac{82\: W*60\: s}{25\: kg*9.8\: N/kg*1.0\: m}=20"

Answer: 20